OpenStudy (anonymous):

The measure of angle ABD is represented by 4y+7. The measure of angle DBC is represented by 3y+12. The measure of angle ABC is represented by 10y-5.

OpenStudy (anonymous):

Therefore, when you sub in the values for each angle, you get: ABC = ABD + DBC 10y - 5 = (4y + 7) + (3y + 12) With a little algebra, you get: 10 y - 5 = 7y + 19 3 y = 24 y = 8 Now that you have y, you can go back and substitute into the original expressions given for each angle. ABD = 4y+ 7 = 4 (8) + 7 = 39 DBC = 3y + 12 = 3 (8) + 12 = 36 ABC = 10y - 5 = 10 (8) - 5 = 75. You can do a check as well, to make sure it works. Left Side Right Side = ABC = ABD + DBC = 75 = 39 + 36 = 75 = 75 "Left side equals right side, therefore, the solution is correct."

OpenStudy (anonymous):

ok big help.